A recent study found that the average life expectancy of a person living in Africa is 53 years with a standard deviation of 7.5 years. If a person is selected at random, what is the probability that the person will die before the age of 65?
A) 94.52%
B) 82.89%
C) 94.95%
D) 88.49%

Answer by stanbon(48568)

A recent study found that the average life expectancy of a person living in Africa is 53 years with a standard deviation of 7.5 years. If a person is selected at random, what is the probability that the person will die before the age of 65?
-----------
Find the z value of 65 and the z value of 0
z(0) = (0-53)/7.5 = .
z(65) = (65-53)/7.5 = 1.6
P(0 < x < 65) = P(-7.067 < z < 1.6) = 0.9452
A) 94.52%
B) 82.89%
C) 94.95%
D) 88.49%
----------------------------
================
Cheers,
Stan H.